OpenStudy (anonymous):

how do you write the trigonometric expression in terms of sine and cosine and simplify: cos(x)tan(x)

OpenStudy (abb0t):

\[\cos(x)\frac{ \sin(x) }{ \cos(x) }\] simplify to get and you get sin(x)

OpenStudy (anonymous):

oohhhh thank you tons!!!